# On a question of Blecher, Pisier, Shlyakhtenko

### v1

@@ 1. Introduction and main results
where $\gamma_{2}^{*}$ is dual of the $\gamma_{2}$ norm, derived from the Hilbert space factorization norm for linear maps. This theorem highlights a remarkable property of tensor products of Banach spaces, revealing deep connections between different norms and structures (see [5] and the references therein).
In 1974 Pisier provided a positive answer to one of Grothendieck’s problems by showing that
(1.2) $A^{*}\otimes_{\epsilon}B^{*}\cong A^{*}\otimes_{\gamma_{2}^{*}}B^{*}\quad\mbox{isomorphically}$
holds for $C^{*}$-algebra $A$, $B$ with an additional approximability assumption [16]. That is, Pisier replaced commutative $L_{1}$-spaces by noncommutative ones. This result was later improved by Haagerup [9], who removed the approximability assumption made in [16] and clarified the correct analogue to $\gamma_{2}^{*}$ in the context of operator algebras.
@@
where $A^{*}\otimes_{\mu}B^{*}=A^{*}\otimes_{h}B^{*}+B^{*}\otimes_{h}A^{*}$ may be considered as a quotient of the direct sum $A^{*}\otimes_{h}B^{*}\oplus B^{*}\otimes_{h}A^{*}$ with respect to the map $q(a_{1}\otimes b_{1}\oplus b_{2}\otimes a_{2})=a_{1}\otimes b_{1}+a_{2}\otimes b_{2}$.
A breakthrough in this problem was achieved by Piser and Shlyakhtenko in [20], where they conclusively demonstrated that the answer to the preceding question is indeed positive at the Banach space level. They proved
@@
assuming that the $C^{*}$-algebras $A$ and $B$ are exact. Notably, their results extended beyond the setting of $C^{*}$-algebras to encompass general exact operator spaces. Subsequently, Haagerup and Musat in [10] furthered this understanding to include general $C^{*}$-algebras. However, the broader question concerning the operator space structure of the corresponding tensors products in Conjecture (1.3) remained unresolved in [20] (see also [18, Problem 21.2]).
@@
$C\otimes_{\min}(\mathcal{K}(\ell_{2})^{*}\otimes_{\min}\mathcal{K}(\ell_{2})^{*})\neq C\otimes_{\min}(\mathcal{K}(\ell_{2})^{*}\otimes_{\mu}\mathcal{K}(\ell_{2})^{*})\quad\text{isomorphically}\,,$
@@
Since it is well known that $C^{*}\langle\ell_{1}\rangle=C^{*}(\mathbb{F}_{\infty})$, one can alternatively express Theorem 1.2 by indicating that the natural inclusion $OA(\ell_{1})\otimes_{max}OA(\ell_{1})\subseteq C^{*}(\mathbb{F}_{\infty})\otimes_{max}C^{*}(\mathbb{F}_{\infty})$ is not completely isomorphic, even when restricted to the linear subspace $\ell_{1}\otimes\ell_{1}$. According to the precise definitions of $OA(\ell_{1})$ and $C^{*}\langle\ell_{1}\rangle$, Theorem 1.2 underscores a significant limitation in lifting specific sets of commuting contractions to commuting unitaries (see Remark 4.1).
@@ 2.2. Some basic results involving the Haagerup tensor norm
Let $(x_{i})_{i}$ and $(y_{j})_{j}$ be two families of contractions in $\mathcal{B}(H)$ such that $[x_{i},y_{j}]=0$ for every $i,j$. If we assume, in addition, the elements $x_{j}$ to be normal, it follows from Fuglede’s Theorem [7] that $[x_{i},y_{j}^{\dagger}]=0$ for every $i,j$. According to Russo-Dye Theorem, we can then write that for every family of matrices $(a_{i,j})_{i,j}\in M_{n}$, it holds that
$\Big{\|}\sum_{i,j}a_{i,j}\otimes x_{i}y_{j}\Big{\|}_{M_{n}(\mathcal{B}(H))}\,\leq\,\sup\Big{\|}\sum_{i,j}a_{ij}\otimes u_{i}v_{j}\Big{\|}_{M_{n}(\mathcal{B}(H))}\leq\Big{\|}\sum_{ij}a_{ij}\otimes e_{i}\otimes e_{j}\Big{\|}_{M_{n}(\ell_{1}\otimes_{\max}\ell_{1})},\,$
where the supremum runs over all families of commuting unitaries $(u_{i})_{i}$ ,$(v_{j})_{j}$ in $\mathcal{B}(H)$. While Ricard demonstrated [19, Proposition 13.10] the existence of families of commuting contractions $(x_{i})_{i}$ ,$(y_{j})_{j}$ satisfying
$\Big{\|}\sum_{i,j}a_{i,j}\otimes x_{i}y_{j}\Big{\|}_{M_{n}(\mathcal{B}(H))}>\Big{\|}\sum_{ij}a_{ij}\otimes e_{i}\otimes e_{j}\Big{\|}_{M_{n}(\ell_{1}\otimes_{\max}\ell_{1})},\,$
even for scalar coefficients $(a_{i,j})_{i,j}\in\mathbb{C}$ for every $i,j$, Theorem 1.2 provides families of commuting contractions that dramatically violate this inequality - such that even an equivalence between the norms is not possible. In this sense, Theorem 1.2 can be understood as a qualitative witness to the fact that these commuting contractions cannot be normal. We are grateful to Mikael $\o$rdam for reminding us of Fuglede’s result.

### v2

@@ 1. Introduction and main results
where $\gamma_{2}^{*}$ is dual norm of the Hilbert space factorization norm $\gamma_{2}$. This theorem highlights a remarkable property of tensor products of Banach spaces, revealing deep connections between different norms and structures (see [5] and the references therein). In particular, Grothendieck’s theorem implies that every linear map $T:\ell_{\infty}\rightarrow\ell_{1}$ factors through a Hilbert space.
In 1974, Pisier [16] provided a positive answer to one of Grothendieck’s conjectures by showing that
(1.2) $A^{*}\otimes_{\epsilon}B^{*}\cong A^{*}\otimes_{\tilde{\gamma}^{*}_{2}}B^{*}\quad\mbox{isomorphically}$
holds for general $C^{*}$-algebras $A$ and $B$. Here, $\tilde{\gamma}^{*}_{2}$ is the analog to the $\gamma_{2}^{*}$ norm in the context of operator algebras. That is, Pisier replaced commutative $L_{1}$-spaces by noncommutative ones. In particular, Pisier proved that every linear map $T:A\rightarrow B^{*}$ which is approximable, in a particular way, by finite ranks maps, factors through a Hilbert space. This result was later extended by Haagerup [9], who removed such an approximability assumption.
@@
where $A^{*}\otimes_{\mu}B^{*}=A^{*}\otimes_{h}B^{*}+B^{*}\otimes_{h}A^{*}$ may be considered as a quotient of the direct sum $A^{*}\otimes_{h}B^{*}\oplus_{1}B^{*}\otimes_{h}A^{*}$ with respect to the map $q(a_{1}\otimes b_{1}\oplus b_{2}\otimes a_{2})=a_{1}\otimes b_{1}+a_{2}\otimes b_{2}$.
A breakthrough in this problem was achieved by Pisier and Shlyakhtenko in [20], where they conclusively demonstrated that the answer to the preceding question is indeed positive at the Banach space level. They proved
@@
for general $C^{*}$-algebras $A$ and $B$. Notably, their results extended beyond the setting of $C^{*}$-algebras to encompass general exact operator spaces. Subsequently, Haagerup and Musat in [10] furthered the understanding of (1.4), particularly improving the constant involved. However, the broader question concerning the operator space structure of the corresponding tensor products in Conjecture (1.3) remained unresolved in [20] (see also [18, Problem 21.2]).
@@
$C\otimes_{\min}(\mathcal{K}(\ell_{2})^{*}\otimes_{\min}\mathcal{K}(\ell_{2})^{*})\neq C\otimes_{\min}(\mathcal{K}(\ell_{2})^{*}\otimes_{\mu}\mathcal{K}(\ell_{2})^{*})\quad\text{isomorphically}\,.$
@@
Since it is well known that $C^{*}\langle\ell_{1}\rangle=C^{*}(\mathbb{F}_{\infty})$, one can alternatively express Theorem 1.2 by indicating that the natural inclusion $OA(\ell_{1})\otimes_{max}OA(\ell_{1})\subseteq C^{*}(\mathbb{F}_{\infty})\otimes_{max}C^{*}(\mathbb{F}_{\infty})$ is not completely isomorphic, even when restricted to the linear subspace $\ell_{1}\otimes\ell_{1}$. According to the precise definitions of $OA(\ell_{1})$ and $C^{*}\langle\ell_{1}\rangle$, Theorem 1.2 underscores a significant limitation involving commuting contractions and commuting unitaries, or even commuting normal contractions (see Remark 4.1).
@@ 2.2. Some basic results involving the Haagerup tensor norm
Let $(x_{i})_{i}$ and $(y_{j})_{j}$ be two families of contractions in $\mathcal{B}(H)$ such that $[x_{i},y_{j}]=0$ for every $i,j$. If we assume, in addition, the elements $x_{i}$’s to be normal, it follows from Fuglede’s Theorem [7] that $[x_{i},y_{j}^{\dagger}]=0$ for every $i,j$. According to Russo-Dye Theorem (or by just using twice the standard block encoding; that is, given any contraction $T\in\mathcal{B}(H)$, we consider the unitary
$U\,=\,\left(\begin{array}[]{cc}T&-\sqrt{1-TT^{\dagger}}\\ \sqrt{1-T^{\dagger}T}&T^{\dagger}\end{array}\right)\in M_{2}(\mathcal{B}(H))),$
we obtain that for every family of matrices $(a_{ij})_{i,j}\in M_{n}$, it holds that
$\Big{\|}\sum_{i,j}a_{ij}\otimes x_{i}y_{j}\Big{\|}_{M_{n}(\mathcal{B}(H))}\,\leq\,\sup\Big{\|}\sum_{i,j}a_{ij}\otimes u_{i}v_{j}\Big{\|}_{M_{n}(\mathcal{B}(H))}\leq\Big{\|}\sum_{ij}a_{ij}\otimes e_{i}\otimes e_{j}\Big{\|}_{M_{n}(\ell_{1}\otimes_{\max}\ell_{1})},\,$
where the supremum runs over all families of commuting unitaries $(u_{i})_{i}$, $(v_{j})_{j}$ in $\mathcal{B}(H)$. While Ricard demonstrated [19, Proposition 13.10] the existence of families of commuting contractions $(x_{i})_{i}$, $(y_{j})_{j}$ and (even scalar) coefficients $(a_{ij})_{i,j}\in\mathbb{C}$ satisfying
$\Big{\|}\sum_{i,j}a_{ij}\otimes x_{i}y_{j}\Big{\|}_{M_{n}(\mathcal{B}(H))}>\Big{\|}\sum_{ij}a_{ij}\otimes e_{i}\otimes e_{j}\Big{\|}_{M_{n}(\ell_{1}\otimes_{\max}\ell_{1})},\,$
Theorem 1.2 provides families of commuting contractions that dramatically violate this inequality - such that even an equivalence between the norms is not possible. Note that considering matrix coefficients $(a_{ij})_{i,j}$ is crucial here, since Grothendieck’s inequality (1.1) implies that
$\Big{\|}\sum_{i,j}a_{ij}x_{i}y_{j}\Big{\|}_{\mathcal{B}(H)}\leq K_{G}\Big{\|}\sum_{ij}a_{ij}e_{i}\otimes e_{j}\Big{\|}_{\ell_{1}\otimes_{\max}\ell_{1}},\,$
for every families of (non-necessarily commuting) contractions $(x_{i})_{i}$ and $(y_{j})_{j}$. Hence, Theorem 1.2 can be understood as a qualitative witness to the fact that the commuting contractions $(x_{i})_{i}$ and $(y_{j})_{j}$ cannot be normal. We are grateful to Mikael $\o$rdam for reminding us of Fuglede’s result.